Two voltmeters, one having a solution of silver salt and the other of a trivalent-metal salt, are connected in series and a current of 2 A is maintained for 1.50 hours. It is found that 1.00 g of the trivalent-metal is deposited.
a) What is the atomic weight of the trivalent metal?
b) How much silver is deposited during this period? Atomic weight of silver is 1.07.9 g/mole.

Grade:11

1 Answers

Aditi Chauhan
9 years ago
Sol. w base 1 = Zit ⇒ 1 = mm/3*96500 *2*1.5*3600 ⇒ mm = 3 * 96500/2 * 1.5 * 3600 = 26.8 g/mole E base 1/E base 2 = w base 1/w base 2 ⇒ 107.9/(mm/3) = w base 1/1 = 107.9 * 3/26.8 = 12.1 gm